OpenStudy (anonymous):

Find the derivative of the function g(x) = ln|x^2 + 6|

OpenStudy (anonymous):

I don't know how to prove it but it is == 2x/x^2+6

OpenStudy (anonymous):

sorry but i dont think thats correct.

OpenStudy (anonymous):

ok I forgot the rules I can't help you after now I'm sorry

OpenStudy (anonymous):

He is correct, Where did you see another answer?

OpenStudy (anonymous):

Derivative of a log function of x: \[\frac{d}{dx}(\log_af(x)) = \frac{f'(x)}{f(x)\ln(a)}\] in the event of natural log [ie ln function] the base of the log is 'e' and the ln(e) is one so you don't see it in the denominator

OpenStudy (anonymous):